Do the following two lines intersect? Answer yes or no, together with the point, if any
5x+8y=-5
-X-1.6y=14

Answer by Boreal(15235)

5x+8y=-5
-X-1.6y=14
multiply the second equation by 5
-5x-8y=70
add the two
0x+0y=65
they do not intersect.
---------------
write in slope intercept form
8y=-5x-5
y=(-5/8)x-(5/8)
the second
-1.6y=-x+14
y=-1/-1.6 +14/-1.6
but 1/1.6 is 10/16 or 5/8
y=(5/8)x-8 2/3
The two lines have the same coefficient in front of the x, and the intercepts are different. They are parallel lines by definition.